OpenStudy (anonymous):

Sal's Sandwich Shop sells wraps and sandwiches as part of its lunch specials. The profit on every sandwich is $2 and the profit on every wrap is $3. Sal made a profit of $1,470 from lunch specials last month. The equation 2x + 3y = 1,470 represents Sal's profits last month, where x is the number of sandwich lunch specials sold and y is the number of wrap lunch specials sold. Change the equation to slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all your work. Describe how you would graph this line using the slope-intercept method.

OpenStudy (anonymous):

4 Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. 5 Graph the function. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work or you may use graphing technology. 6 Suppose Sal's total profit on lunch specials for the next month is $1,593. The profit amounts are the same: $2 for each sandwich and $3 for each wrap. In a paragraph of at least three complete sentences, explain how the graphs of the functions for the two months are similar and how they are different. 7 Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations.

OpenStudy (jadatamiamaraj):

1. Change the equation to slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all your work. 2x + 3y = 1,470 –2x –2x 3y = –2x + 1,470 3y/ 3 = –2x/3 + 1,470/3 y = –2x/3 + 490 the y-intercept of this equation is 490 and the slope is -2x/3. 2. Describe how you would graph this line using the slope-intercept method. Be sure to write using complete sentences. I would graph this line by finding the y-intercept 490 first. Then, you would move down twice from 490 to 488. Finally, you would move to the right 3 times 3. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. F(x) = –2x/3 + 490 the graph would represent the profit of Sal’s Sandwich Shop on every sandwich and wrap from the lunch specials last month. thats how far i got i dont know if its all right tho...